OpenStudy (anonymous):

Why would the domain of f(x)= sec x-1/x^2 be x does not =0, pi/2+npi? i get 0, but not pi/2+npi

jimthompson5910 (jim_thompson5910):

sec(x) = 1/cos(x) so cos(x) can't be zero either

jimthompson5910 (jim_thompson5910):

cos(x) = 0 is true when x = pi/2, 3pi/2, etc... in general, the solutions are: x = pi/2 + n*pi where n is any integer